OpenStudy (anonymous):

Which of the following systems of equations has no solution? a. 18x − 6y = 14; 3y = 9x − 7 b. 7x − 4 = 3y; 6y − 3 = 14x c. 4x + y = 10; 2y = 8x − 2 d. -5y = x − 6; 3x − 15y = 9

OpenStudy (anonymous):

Try B...

OpenStudy (anonymous):

@matricked : is it B ?

OpenStudy (anonymous):

YUP

OpenStudy (anonymous):

7x - 3y = 4 Multiply it by 2, 14x - 6y = 8 14x - 6y = -3 When you subtract them both x and y will get cancel..

OpenStudy (anonymous):

these two equations represent non overlapping paralel lines hence they do not intersect so no solution